OpenStudy (anonymous):

How do I solve this using difference of squares. I understand (a+b)(a-b) 16(x^2+1)^2-4(2x)^2

OpenStudy (anonymous):

\[16(x^{2}+1)^{2}-4(2x)^{2} = 4^{2}*(x^{2}+1)^{2}-2^{2}*(2x)^{2}\] = \[[4*(x^{2}+1)]^{2} - (4x)^{2}\] so let it be a = \[4*(x^{2}+1)\] and b = \[4x\]

OpenStudy (anonymous):

\[a^{2}-b ^{2}= (a-b)(a+b) = (4x ^{2}+4-4x)(4x ^{2}+4+4x)\]

OpenStudy (anonymous):

cool thanks

OpenStudy (anonymous):

you're welcome.